OpenStudy (anonymous):

Medal & fan Using a directrix of y = -2 and a focus of (1, 6), what quadratic function is created?

OpenStudy (anonymous):

@Nnesha @zepdrix @MeganEdward

OpenStudy (anonymous):

or @dan815 or @merchandize lol

Nnesha (nnesha):

 ;-;

OpenStudy (jhannybean):

Recall that the vertex of a parabola lies halfway between the focus and the directrix

Nnesha (nnesha):

alright let's pick point (x,y)  we need to find distance between the point (x,y) and focus and the distance between point and directirx \[\sqrt{(x - a)^2+(y -b)^2}= \left| y-c \right|\] \[\sqrt{(x- 1)^2 + (y-6)^2} = (y-(-2))^2\]